# A Manifold Perspective on the Statistical Generalization of Graph Neural Networks

### v1

@@ 3.2 Manifold neural networks
We consider a $d$-dimensional compact, smooth and differentiable submanifold ${\mathcal{M}}$ embedded in ${\mathbb{R}}^{\mathsf{M}}$. This induces a probability measure $\mu$ over the manifold with density function $\rho:{\mathcal{M}}\rightarrow(0,\infty)$, assumed to be bounded as $0<\rho_{min}\leq\rho(x)\leq\rho_{max}<\infty$ for all $x\in{\mathcal{M}}$. The manifold data supported on each point $x\in{\mathcal{M}}$ is defined by scalar functions $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$ [34]. We use $L^{2}({\mathcal{M}})$ to denote $L^{2}$ functions over ${\mathcal{M}}$ with respect to measure $\mu$. The manifold with probability density function $\rho$ is equipped with a weighted Laplace operator [36], generalizing the Laplace-Beltrami operator as
@@
The manifold convolution operation is defined relying on the Laplace operator ${\mathcal{L}}_{\rho}$ and on the heat diffusion process over the manifold [34]. For a function $f\in L^{2}({\mathcal{M}})$, the manifold diffusion can be explicitly written as $e^{-{\mathcal{L}}_{\rho}}f$. A manifold convolutional filter [34] can be defined in a diffuse-and-sum manner as
@@
with the $k$-th diffusion scaled with a filter parameter $h_{k}\in{\mathbb{R}}$. We consider the case in which the Laplace operator is self-adjoint and positive-semidefinite and the manifold ${\mathcal{M}}$ is compact, in this case, ${\mathcal{L}}_{\rho}$ has real, positive and discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$, written as ${\mathcal{L}}_{\rho}\bm{\phi}_{i}=\lambda_{i}\bm{\phi}_{i}$ where $\bm{\phi}_{i}$ is the eigenfunction associated with eigenvalue $\lambda_{i}$. The eigenvalues are ordered in increasing order as $0<\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots$, and the eigenfunctions are orthonormal and form an eigenbasis of $L^{2}({\mathcal{M}})$. When mapping a manifold signal onto the eigenbasis $[\hat{f}]_{i}=\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}$, the manifold convolution can be written in the spectral domain as
@@
where $f_{l-1}^{q}$, $1\leq q\leq F_{l-1}$ is the $q$-th input feature from layer $l-1$ and $f_{l}^{p}$, $1\leq p\leq F_{l}$ is the $p$-th output feature of layer $l$. We denote MNN as a mapping $\boldsymbol{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)$, where ${\mathbf{H}}\in{\mathcal{H}}\in{\mathbb{R}}^{P}$ is a collective set of filter parameters in all the manifold convolutional filters.
@@ 4 Generalization analysis of GNNs based on manifold
We consider a manifold ${\mathcal{M}}$ as defined in Section 3.2, equipped with a weighted Laplace operator ${\mathcal{L}}_{\rho}$ as (4). Since functions $f\in L^{2}({\mathcal{M}})$ are used to characterize information over manifold ${\mathcal{M}}$, we restrict our analysis to a finite-dimensional subset of $L^{2}({\mathcal{M}})$ up to some eigenvalue of ${\mathcal{L}}_{\rho}$, which is defined as a bandlimited signal.
@@ 4.1 Manifold label prediction via node label prediction
$\left|\hat{h}(\lambda)\right|={\mathcal{O}}\left(\lambda^{-d}\right),\quad\lambda\in(0,\infty).$ (11)
@@
$GA=\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})-\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathbf{G}}({\mathbf{H}}).$ (14)
@@
Theorem 1 shows that the generalization gap decreases approximately linearly with the number of nodes $N$ in the logarithmic scale. It also increases with the dimension of the underlying manifold $d$. Another observation is that the generalization gap scales with the size of GNN architectures. Specifically, it scales polynomially with the number of features $F$, while the exponent depends on the number of layers $L$. This generalization result attests that when the graph is supported on a manifold, a GNN trained on this discrete graph model can approximate the MNN performance. GNNs trained on these finite sampled points can generalize to predict the labels of other unseen points over the manifold with large enough nodes in the graph. To generalize effectively over a manifold with a higher dimension, more sampled points are required to construct the graph to guarantee the generalization capability.
@@ 4.2 Manifold classification via graph classification
$GA=\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathcal{M}}({\mathbf{H}})-\min_{{\mathbf{H}}\in{\mathcal{H}}}R_{\mathbf{G}}({\mathbf{H}}).$ (18)
@@ Appendix A Induced manifold signals
Let $\{V_{i}\}_{i=1}^{N}$ be the Voronoi decomposition [48] of ${\mathcal{M}}$ with respect to $X_{N}$ with $V_{i}\subset B_{r}(x_{i})$ and $\text{vol}(V_{i})=\mu_{i}=1/N$, where $B_{r}(x_{i})$ denoted the closed metric ball of radius $r$ centered at $x_{i}\in{\mathcal{M}}$. The Voronoi decomposition can be explicitly denoted as
$V_{i}=\{x\in{\mathcal{M}}:\text{dist}(x_{i},x)\leq\text{dist}(x_{j},x),j=1,2\cdots i-1,i+1,\cdots N\},$ (21)
where dist represents the curvature distance between two points on the manifold [37]. The radius of $V_{i}$ can be bounded as $r\sim(\log N/N)^{1/d}$ [49].
@@ Appendix B Convergence of GNN to MNN
The convergence of GNN on sampled graphs to MNN provides the support for the generalization analysis.
@@
$\frac{1}{N}\sum_{i=1}^{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}}^{*},{\mathbf{L}}_{N},{\mathbf{x}})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}}^{*},{\mathcal{L}}_{\rho},{\mathbf{I}}_{N}{\mathbf{x}})\|_{2}\leq LF^{L-1}\left(C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}}\right)$ (23)
@@
with $a_{i}\in\{-1,1\}$ for all $i<K$ and $\theta$ the eigengap of ${\mathcal{L}}$, i.e., $\theta_{i}=\min\{\lambda_{i}-\lambda_{i-1},\lambda_{i+1}-\lambda_{i}\}$. The constants $C_{{\mathcal{M}},1}$, $C_{{\mathcal{M}},2}$ depend on $d$ and the volume of ${\mathcal{M}}$.
@@
The inner product of signals $f,g\in L^{2}({\mathcal{M}})$ is defined as
$\|f\|^{2}_{{\mathcal{M}}}={\langle f,f\rangle_{{\mathcal{M}}}}.$ (26)
@@
$\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|=O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right).$ (27)
@@
$\|{\mathbf{h}}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f-{\mathbf{P}}_{N}{\mathbf{h}}({\mathcal{L}}_{\rho})f\|\leq\Bigg{\|}\sum_{i=1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\Bigg{\|}$ (29)
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}\bm{\phi}_{i}+\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\Bigg{\|}$ (30)
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}\bm{\phi}_{i}\Bigg{\|}+\left\|\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\right\|$ (31)
@@
$\leq\|{\mathbf{P}}_{N}f\|\sum_{i=1}^{M}C_{{\mathcal{M}},1}\epsilon\lambda_{i}^{-d}\leq\|{\mathbf{P}}_{N}f\|C_{{\mathcal{M}},1}\epsilon\sum_{i=1}^{M}i^{-2}$ (34)
@@
In (34), we implement Weyl’s law [51] which indicates that eigenvalues of Laplace operator scales with the order $\lambda_{i}\sim i^{2/d}$. The last inequality comes from the fact that $\sum_{i=1}^{\infty}i^{-2}=\frac{\pi^{2}}{6}$. The second term in (32) can be bounded with the triangle inequality as
@@
$\leq\sum_{i=1}^{M}(\lambda_{i}^{-d})\sqrt{\frac{\log(1/\delta)}{N}}\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (42)
$\leq\frac{\pi^{2}}{6}\sqrt{\frac{\log(1/\delta)}{N}}\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right):=A_{3}(N)$ (43)
@@
with $C_{1}^{\prime}=C_{{\mathcal{M}},1}\frac{\pi^{2}}{6}\|f\|_{\mathcal{M}}$, $C_{2}^{\prime}=C_{{\mathcal{M}},2}\frac{\pi^{2}}{6}$, $C_{3}^{\prime}=\frac{\pi^{2}}{6}$ and $C_{4}^{\prime}=\|f\|_{\mathcal{M}}$. As $N$ goes to infinity, for every $\delta>0$, there exists some $M_{0}$, such that for all $M>M_{0}$ it holds that $A_{4}(M)\leq\delta/2$. There also exists $n_{0}$, such that for all $N>n_{0}$, it holds that $A_{1}(N)+A_{2}(M_{0},N)+A_{3}(N)\leq\delta/2$. We can conclude that the summations converge as $N$ goes to infinity. We see $M$ large enough to have $M^{-1}\leq\delta^{\prime}$, which makes the eigengap $\theta_{M}$ also bounded by some constant. We combine the first two terms as

### v2

@@ 3.2 Manifold neural networks
We consider a $d$-dimensional compact, smooth and differentiable Riemannian submanifold ${\mathcal{M}}$ embedded in ${\mathbb{R}}^{\mathsf{M}}$ with finite volume. This induces a Hausdorff probability measure $\mu$ over the manifold with density function $\rho:{\mathcal{M}}\rightarrow(0,\infty)$, assumed to be bounded as $0<\rho_{min}\leq\rho(x)\leq\rho_{max}<\infty$ for all $x\in{\mathcal{M}}$. The manifold data supported on each point $x\in{\mathcal{M}}$ is defined by scalar functions $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$ [34]. We use $L^{2}({\mathcal{M}})$ to denote $L^{2}$ functions over ${\mathcal{M}}$ with respect to measure $\mu$. The manifold with probability density function $\rho$ is equipped with a weighted Laplace operator [36], generalizing the Laplace-Beltrami operator as
@@
The manifold convolution operation is defined relying on the Laplace operator ${\mathcal{L}}_{\rho}$ and on the heat diffusion process over the manifold [34]. For a function $f\in L^{2}({\mathcal{M}})$ as the initial heat condition over ${\mathcal{M}}$, the heat condition diffused by a unit time step can be explicitly written as $e^{-{\mathcal{L}}_{\rho}}f$. A manifold convolutional filter [34] can be defined in a diffuse-and-sum manner as
@@
with the $k$-th diffusion scaled with a filter parameter $h_{k}\in{\mathbb{R}}$. We consider the case in which the Laplace operator is self-adjoint with respect to the inner product defined in (24) and positive-semidefinite and the manifold ${\mathcal{M}}$ is compact, in this case, ${\mathcal{L}}_{\rho}$ has real, positive and discrete eigenvalues $\{\lambda_{i}\}_{i=1}^{\infty}$, written as ${\mathcal{L}}_{\rho}\bm{\phi}_{i}=\lambda_{i}\bm{\phi}_{i}$ where $\bm{\phi}_{i}$ is the eigenfunction associated with eigenvalue $\lambda_{i}$. The eigenvalues are ordered in increasing order as $0=\lambda_{1}\leq\lambda_{2}\leq\lambda_{3}\leq\ldots$, and the eigenfunctions are orthonormal and form an eigenbasis of $L^{2}({\mathcal{M}})$. When mapping a manifold signal onto the eigenbasis $[\hat{f}]_{i}=\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}$, the manifold convolution can be written in the spectral domain as
@@
where $f_{l-1}^{q}$, $1\leq q\leq F_{l-1}$ is the $q$-th input feature from layer $l-1$ and $f_{l}^{p}$, $1\leq p\leq F_{l}$ is the $p$-th output feature of layer $l$. We denote MNN as a mapping $\boldsymbol{\Phi}({\mathbf{H}},{\mathcal{L}}_{\rho},f)$, where ${\mathbf{H}}\in{\mathcal{H}}\subset{\mathbb{R}}^{P}$ is a collective set of filter parameters in all the manifold convolutional filters.
@@ 4 Generalization analysis of GNNs based on manifold
We consider a manifold ${\mathcal{M}}$ as defined in Section 3.2, with a weighted Laplace operator ${\mathcal{L}}_{\rho}$ as (4). Since functions $f\in L^{2}({\mathcal{M}})$ characterize information over manifold ${\mathcal{M}}$, we restrict our analysis to a finite-dimensional subset of $L^{2}({\mathcal{M}})$ up to some eigenvalue of ${\mathcal{L}}_{\rho}$, defined as a bandlimited signal.
@@ 4.1 Manifold label prediction via node label prediction
$\left|\hat{h}(\lambda)\right|={\mathcal{O}}\left(\lambda^{-d}\right),\quad\left|\hat{h}^{\prime}(\lambda)\right|\leq C_{L}\lambda^{-d-1},\quad\lambda\in(0,\infty).$ (11)
@@
$GA=\sup_{{\mathbf{H}}\in{\mathcal{H}}}\left|R_{\mathcal{M}}({\mathbf{H}})-R_{\mathbf{G}}({\mathbf{H}})\right|.$ (14)
@@
Theorem 1 shows that the generalization gap decreases approximately linearly with the number of nodes $N$ in the logarithmic scale. It also increases with the dimension of the underlying manifold $d$. Another observation is that the generalization gap scales with the size of GNN architectures. Specifically, it scales polynomially with the number of features $F$, while the exponent depends on the number of layers $L$. This generalization result attests that when the graph is supported on a manifold, a GNN trained on this discrete graph model can approximate the MNN performance. GNNs trained on these finite sampled points can generalize to predict the labels of other unseen points over the manifold with large enough nodes in the graph. To generalize effectively over a manifold with a higher dimension, we need more sampled points to construct the graph to guarantee the generalization.
@@ 4.2 Manifold classification via graph classification
$GA=\sup_{{\mathbf{H}}\in{\mathcal{H}}}\left|R_{\mathcal{M}}({\mathbf{H}})-R_{\mathbf{G}}({\mathbf{H}})\right|.$ (18)
@@ Appendix A Induced manifold signals
Let $\mu_{N}$ be the empirical measure of the random sample as
$\mu_{N}=\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}.$ (21)
Let $\{V_{i}\}_{i=1}^{N}$ be the decomposition [48] of ${\mathcal{M}}$ with respect to $X_{N}$ with $V_{i}\subset B_{r}(x_{i})$, where $B_{r}(x_{i})$ denoted the closed metric ball of radius $r$ centered at $x_{i}\in{\mathcal{M}}$. The decomposition can be achieved by the optimal transportation map $T:{\mathcal{M}}\rightarrow X_{N}$, which is defined by the $\infty$-Optimal Transport distance between $\mu$ and $\mu_{N}$.
$d_{\infty}(\mu,\mu_{N}):=\min_{T:T_{\#}\mu=\mu_{N}}\text{esssup}_{x\in{\mathcal{M}}}d(x,T(x)),$ (22)
where $T_{\#}\mu=\mu_{N}$ indicates that $\mu(T^{-1}(V))=\mu_{N}(V)$ for every $V_{i}$ of ${\mathcal{M}}$. This transportation map $T$ induces the partition $V_{1},V_{2},\cdots V_{N}$ of ${\mathcal{M}}$, where $V_{i}:=T^{-1}(\{x_{i}\})$ with $\mu(V_{i})=\frac{1}{N}$ for all $i=1,\cdots N$. The radius of $V_{i}$ can be bounded as $r\leq A(\log N/N)^{1/d}$ with $A$ related to the geometry of ${\mathcal{M}}$ [48, Theorem 2].
@@ Appendix B Convergence of GNN to MNN
The convergence of GNN on sampled graphs to MNN provides the support for the generalization analysis. We first introduce the inner product over the manifold. The inner product of signals $f,g\in L^{2}({\mathcal{M}})$ is defined as
$\|f\|^{2}_{{\mathcal{M}}}={\langle f,f\rangle_{{\mathcal{M}}}}.$ (25)
@@
$\frac{1}{N}\sum_{i=1}^{N}\|\bm{\Phi}_{\mathbf{G}}({\mathbf{H}},{\mathbf{L}}_{N},{\mathbf{x}})-{\mathbf{P}}_{N}{\bm{\Phi}}({\mathbf{H}},{\mathcal{L}}_{\rho},{\mathbf{I}}_{N}{\mathbf{x}})\|_{2}\leq LF^{L-1}\left(C_{1}\epsilon+C_{2}\sqrt{\frac{\log(1/\delta)}{N}}\right)$ (26)
@@
with $a_{i}\in\{-1,1\}$ for all $i<K$ and $\theta$ the eigengap of ${\mathcal{L}}$, i.e., $\theta_{i}=\min\{\lambda_{i}-\lambda_{i-1},\lambda_{i+1}-\lambda_{i}\}$. The constants $C_{{\mathcal{M}},1}$, $C_{{\mathcal{M}},2}$ depend on $d$ and the volume, the injectivity radius and sectional curvature of ${\mathcal{M}}$.
@@
$\left|\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|=O\left(\sqrt{\frac{\log(1/\delta)}{N}}\right).$ (28)
@@
$\|{\mathbf{h}}({\mathbf{L}}_{N}){\mathbf{P}}_{N}f-{\mathbf{P}}_{N}{\mathbf{h}}({\mathcal{L}}_{\rho})f\|=\Bigg{\|}\sum_{i=1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\Bigg{\|}$ (30)
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}+\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\Bigg{\|}$ (31)
$\leq\Bigg{\|}\sum_{i=1}^{M}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}-\sum_{i=1}^{M}\hat{h}(\lambda_{i})\langle f,\bm{\phi}_{i}\rangle_{{\mathcal{M}}}{\mathbf{P}}_{N}\bm{\phi}_{i}\Bigg{\|}+\left\|\sum_{i=M+1}^{N}\hat{h}(\lambda_{i,N})\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle\bm{\phi}_{i,N}\right\|$ (32)
@@
$\leq\|{\mathbf{P}}_{N}f\|\sum_{i=1}^{M}|\hat{h}^{\prime}(\lambda_{i})||\lambda_{i,N}-\lambda_{i}|$ (35)
$\leq\|{\mathbf{P}}_{N}f\|\sum_{i=1}^{M}C_{{\mathcal{M}},1}C_{L}\epsilon\lambda_{i}^{-d}$ (36)
$\leq\|{\mathbf{P}}_{N}f\|C_{L}C_{{\mathcal{M}},1}\epsilon\sum_{i=1}^{M}i^{-2}$ (37)
@@
In (36), it depends on the low-pass filter assumption in Definition 4. In (37), we implement Weyl’s law [50] which indicates that eigenvalues of Laplace operator scales with the order $\lambda_{i}\sim i^{2/d}$. The last inequality comes from the fact that $\sum_{i=1}^{\infty}i^{-2}=\frac{\pi^{2}}{6}$. The second term in (33) can be bounded with the triangle inequality as
@@
$\leq\sum_{i=1}^{M}(\lambda_{i,N}^{-d})\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (45)
$\leq\sum_{i=1}^{M}(1+C_{{\mathcal{M}},1}\epsilon)^{-d}(\lambda_{i}^{-d})\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (46)
$\leq\frac{\pi^{2}}{6}\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right):=A_{3}(N)$ (47)
The term $\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ can be decomposed by inserting a term $\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle$ as
$\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ $\leq\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle+\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ (48)
$\leq\left|\langle{\mathbf{P}}_{N}f,\bm{\phi}_{i,N}\rangle-\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle\right|+\left|\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ (49)
$\leq\|{\mathbf{P}}_{N}f\|\|\bm{\phi}_{i,N}-{\mathbf{P}}_{N}\bm{\phi}_{i}\|+\left|\langle{\mathbf{P}}_{N}f,{\mathbf{P}}_{N}\bm{\phi}_{i}\rangle-\langle f,\bm{\phi}_{i}\rangle_{\mathcal{M}}\right|$ (50)
$\leq\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)\frac{C_{{\mathcal{M}},2}\lambda_{i}\epsilon}{\theta_{i}}+\sqrt{\frac{\log(1/\delta)}{N}}$ (51)
Then equation (46) can be bounded as
$\leq\sum_{i=1}^{M}(1+C_{{\mathcal{M}},1}\epsilon)^{-d}(\lambda_{i}^{-d})\left(\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)\frac{C_{{\mathcal{M}},2}\lambda_{i}\epsilon}{\theta_{i}}+\sqrt{\frac{\log(1/\delta)}{N}}\right)\left(1+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)$ (52)
$\leq\frac{\pi^{2}}{6}\max_{i=1,\cdots,M}\frac{C_{{\mathcal{M}},2}\epsilon}{\theta_{i}}\left(\|f\|_{\mathcal{M}}+\left(\frac{\log(1/\delta)}{N}\right)^{\frac{1}{4}}\right)+\frac{\pi^{2}}{6}\sqrt{\frac{\log(1/\delta)}{N}}$ (53)
@@
with $C_{1}^{\prime}=C_{L}C_{{\mathcal{M}},1}\frac{\pi^{2}}{6}\|f\|_{\mathcal{M}}$, $C_{2}^{\prime}=C_{{\mathcal{M}},2}\frac{\pi^{2}}{6}$, $C_{3}^{\prime}=\frac{\pi^{2}}{6}$ and $C_{4}^{\prime}=\|f\|_{\mathcal{M}}$. As $N$ goes to infinity, for every $\delta>0$, there exists some $M_{0}$, such that for all $M>M_{0}$ it holds that $A_{4}(M)\leq\delta/2$. There also exists $n_{0}$, such that for all $N>n_{0}$, it holds that $A_{1}(N)+A_{2}(M_{0},N)+A_{3}(N)\leq\delta/2$. We can conclude that the summations converge as $N$ goes to infinity. We see $M$ large enough to have $M^{-1}\leq\delta^{\prime}$, which makes the eigengap $\theta_{M}$ also bounded by some constant. We combine the first two terms as